# VITA: Variational Pretraining of Transformers for Climate-Robust Crop Yield Forecasting

### v2

@@ Conclusion
We introduced VITA, a variational pretraining framework for forecasting crop yield under extreme weather. By leveraging satellite weather datasets through decoder-free variational learning and seasonality-aware priors, VITA achieves state-of-the-art performance on extreme weather years (R² = 0.729 for corn, 0.727 for soybeans). The framework requires only basic weather variables available globally, allowing deployment in data-scarce regions where multi-modal approaches are infeasible. Our comprehensive evaluation across 763 US Corn Belt counties with statistical validation and spatial transfer experiments, demonstrates robust performance, and provides a scalable approach for AI-driven food security.
@@
where $\hat{y_{j}}=\mathbb{E}_{q_{\phi}(z_{j}\mid x_{j})}[\mu_{\theta}(z_{j},y_{\text{past},j})]$ is the predicted yield. Note that Equation (20) is still the variational lower bound and the $\beta$ multiplyer does not originate from the weakening of the variational objective.
@@
Our approach leverages two distinct spatial datasets with different coverage and resolution. The pretraining dataset contains 116 spatial grids at 0.5° resolution covering the continental United States, Central America and South America, with 108 grids used for training and 8 grids for validation. The fine-tuning dataset focuses on 763 counties within the US Corn Belt region, representing the core agricultural areas for corn and soybean production. Figure 7 illustrates the spatial coverage of both datasets, demonstrating how weather knowledge learned from the broader Americas region transfers effectively to US agricultural counties.
@@
Table 14 shows a list of 31 weather measurements in the pretraining data. The first 28 measurements were downloaded from the NASA Power Project (NASA 2024) from 1984 to 2022. The last three were predicted using Tetens equation (Vapor Pressure and Vapor Pressure Deficit) and FAO-Penman-Monteith equation (Reference Evapotranspiration). The data was downloaded for daily measurements and weekly mean was computed for each variable.
@@
We pretrain on weather sequences using batch size 256, learning rate $5\times 10^{-4}$, and maximum context length of 364 weeks (7 years). Progressive masking increases from 10 to 25 features over 100 epochs, with 10-epoch linear warmup followed by exponential decay ($\delta=0.99$). The variational objective uses $\alpha=0.5$ to balance reconstruction and KL regularization.
@@
Table 7 details the VITA architecture. The core transformer encoder uses 4 layers with 10 attention heads, hidden dimension $d_{h}=200$, and MLP dimension $d_{\text{mlp}}=800$, processing 364-week sequences. The weather output projection produces mean and variance ($31\times 2$) for the variational posterior. Weather attention aggregates 31 feature dimensions using a learned attention mechanism (hidden dim 16). The sinusoidal prior parameterizes amplitude, frequency, phase, and variance for 31 weather variables across 364 time positions ($124\times 364$ total parameters). All MLPs use GELU activation (Hendrycks and Gimpel 2017).
@@
Pretraining on 4×L40S GPUs takes 27-29 minutes, extrapolating to  108-116 minutes on a single L40S. Combined with 25-minute fine-tuning, the complete pipeline requires  133-141 minutes (¡2.5 hours) on a single GPU. The full hyperparameter grid search (27 configurations × 2 crops = 54 experiments) totals  22.5-23.4 hours of fine-tuning time.
@@
Figure 6: Hyperparameter robustness on extreme years: VITA-Sinusoidal consistently outperforms other baselines across all configurations tested. The CNN-RNN performs considerably worse and not shown in the plot.
@@
Crop Model Pretrained Best R² Mean R² t-statistic p-value
Corn VITA-Std No 0.667 $0.463\pm 0.230$ 3.82 $7.5\times 10^{-4}$
Non-US 0.662 $0.632\pm 0.018$
VITA-Sinusoidal No 0.672 $0.469\pm 0.227$ 3.61 $1.3\times 10^{-3}$
Non-US 0.658 $0.627\pm 0.020$
Soybean VITA-Std No 0.689 $0.575\pm 0.134$ 3.72 $9.5\times 10^{-4}$
Non-US 0.710 $0.674\pm 0.020$
VITA-Sinusoidal No 0.684 $0.569\pm 0.139$ 3.685 $1.0\times 10^{-3}$
Non-US 0.707 $0.669\pm 0.021$
Table 10: Spatial transfer learning: pretraining on Central/South America (excluding continental US) across 27 hyperparameter configurations. P-values from paired t-tests comparing Non-US pretrained vs no pretraining.
Tables 9 and 10 comprehensively evaluate pretraining benefits across hyperparameter configurations (seed 1234). T-BERT shows consistent but moderate gains: +10.8% mean corn R² (0.647 vs 0.584, $t=9.13$, $p<0.001$) and +5.0% soybean R² (0.693 vs 0.660, $t=4.36$, $p<0.01$), with low variance (±0.014-0.020) indicating stable, initialization-insensitive optimization.
@@
To establish a domain-grounded comparison, we implement a linear regression baseline following the USDA Economic Research Service (ERS) yield modeling framework (Westcott and Jewison 2013). This approach uses agronomically-motivated features (temperature and precipitation during critical growth periods) conceptually similar to the NASS Yield Forecasting Program (NASS 2023), which relies on survey-based regression estimates.
@@
Lead Time Performance
Reduced Temporal Context
Crop VITA Prior 7 Years 5 Years
R² (RMSE) R² (RMSE)
Corn Std. Normal 0.706 ± 0.025 (17.1 ± 0.7) 0.666 (18.2)
Sinusoidal 0.729 ± 0.008 (16.3 ± 0.2) 0.697 (17.4)
Soybean Std. Normal 0.698 ± 0.020 (5.2 ± 0.2) 0.694 (5.3)
Sinusoidal 0.722 ± 0.005 (5.0 ± 0.1) 0.684 (5.3)

### v3

@@ Conclusion
We introduced VITA, a variational pretraining framework for forecasting crop yield under extreme weather. It is pretrained on satellite weather datasets through decoder-free variational learning and achieves state-of-the-art performance on extreme weather years (R² = 0.729 for corn, 0.722 for soybeans). VITA requires only basic weather variables available globally, allowing deployment in data-scarce regions where multi-modal approaches are infeasible. Our comprehensive evaluation across 763 U.S. Corn Belt counties with statistical validation and spatial transfer experiments demonstrates robust, real performance.
Reproducibility Checklist
Instructions for Authors:
This document outlines key aspects for assessing reproducibility. Please provide your input by editing this .tex file directly.
For each question (that applies), replace the “Type your response here” text with your answer.
Example: If a question appears as
[…]
Type your response here
you would change it to:
[…]
yes
Please make sure to:
  • Replace ONLY the “Type your response here” text and nothing else.
  • Use one of the options listed for that question (e.g., yes, no, partial, or NA).
[…]
1. General Paper Structure
  • $1.$
    Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes/partial/no/NA) Yes
  • 1.2.
    Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes/no) Yes
  • 1.3.
    Provides well-marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper (yes/no) Yes
2. Theoretical Contributions
  • 2.1.
    Does this paper make theoretical contributions? (yes/no) Yes
    If yes, please address the following points:
    • $2.$
      All assumptions and restrictions are stated clearly and formally (yes/partial/no) Yes
    • 2.3.
      All novel claims are stated formally (e.g., in theorem statements) (yes/partial/no) Yes
    • 2.4.
      Proofs of all novel claims are included (yes/partial/no) Yes
    • 2.5.
      Proof sketches or intuitions are given for complex and/or novel results (yes/partial/no) Yes
    • 2.6.
      Appropriate citations to theoretical tools used are given (yes/partial/no) Yes
    • 2.7.
      All theoretical claims are demonstrated empirically to hold (yes/partial/no/NA) Yes
    • 2.8.
      All experimental code used to eliminate or disprove claims is included (yes/no/NA) NA
3. Dataset Usage
  • 3.1.
    Does this paper rely on one or more datasets? (yes/no) Yes
    If yes, please address the following points:
    • 3.2.
      A motivation is given for why the experiments are conducted on the selected datasets (yes/partial/no/NA) Yes
    • $3.$
      All novel datasets introduced in this paper are included in a data appendix (yes/partial/no/NA) Yes
    • 3.4.
      All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes (yes/partial/no/NA) Yes
    • 3.5.
      All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations (yes/no/NA) Yes
    • 3.6.
      All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available (yes/partial/no/NA) Yes
    • 3.7.
      All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing (yes/partial/no/NA) NA
4. Computational Experiments
  • 4.1.
    Does this paper include computational experiments? (yes/no) Yes
    If yes, please address the following points:
    • 4.2.
      This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting (yes/partial/no/NA) Yes
    • 4.3.
      Any code required for pre-processing data is included in the appendix (yes/partial/no) Yes
    • $4.$
      All source code required for conducting and analyzing the experiments is included in a code appendix (yes/partial/no) Yes
    • 4.5.
      All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes (yes/partial/no) Yes
    • 4.6.
      All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes/partial/no) Yes
    • 4.7.
      If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results (yes/partial/no/NA) Yes
    • 4.8.
      This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks (yes/partial/no) Yes
    • 4.9.
      This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics (yes/partial/no) Yes
    • 4.10.
      This paper states the number of algorithm runs used to compute each reported result (yes/no) Yes
    • 4.11.
      Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information (yes/no) Yes
    • 4.12.
      The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank) (yes/partial/no) Yes
    • 4.13.
      This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments (yes/partial/no/NA) Yes
@@ 4. Computational Experiments
where $\hat{y_{j}}=\mathbb{E}_{q_{\phi}(z_{j}\mid x_{j})}[\mu_{\theta}(z_{j},y_{\text{past},j})]$ is the predicted yield. Note that Equation (20) is still the variational lower bound and the $\beta$ multiplier does not originate from the weakening of the variational objective.
@@
Our approach leverages two distinct spatial datasets with different coverage and resolution. The pretraining dataset contains 116 spatial grids at 0.5° resolution covering the continental United States, Central America and South America, with 108 grids used for training and 8 grids for validation. The fine-tuning dataset focuses on 763 counties within the U.S. Corn Belt region, representing the core agricultural areas for corn and soybean production. Figure 7 illustrates the spatial coverage of both datasets, demonstrating how weather knowledge learned from the broader Americas region transfers effectively to US agricultural counties.
@@
Table 12 shows a list of 31 weather measurements in the pretraining data. The first 28 measurements were downloaded from the NASA Power Project (NASA 2024) from 1984 to 2022. The last three were predicted using Tetens equation (Vapor Pressure and Vapor Pressure Deficit) and FAO-Penman-Monteith equation (Reference Evapotranspiration). The data was downloaded as daily measurements and weekly mean was computed for each variable.
@@
We pretrain on weather sequences using batch size 256, and learning rate $5\times 10^{-4}$. Progressive masking increases from 10 to 25 features over 100 epochs, with 10-epoch linear warmup followed by exponential decay ($\delta=0.99$). The variational objective uses $\alpha=0.5$ to balance reconstruction and KL regularization.
@@
Table 7 details the VITA architecture. The core transformer encoder uses 4 layers with 10 attention heads, hidden dimension $d_{h}=200$, and MLP dimension $d_{\text{mlp}}=800$, processing weather sequences. The weather output projection produces mean and variance ($31\times 2$) for the variational posterior. Weather attention aggregates 31 feature dimensions using a learned attention mechanism (hidden dim 16). The sinusoidal prior parameterizes amplitude, frequency, phase, and variance for 31 weather variables across 364 time positions ($124\times 364$ total parameters). All MLPs use GELU activation (Hendrycks and Gimpel 2017).
@@
Pretraining on 4×L40S GPUs takes 27-29 minutes, extrapolating to  108-116 minutes on a single L40S. Combined with 25-minute fine-tuning, the complete pipeline requires  133-141 minutes (<2.5<2.5 hours) on a single GPU. The full hyperparameter grid search (27 configurations × 2 crops = 54 experiments) totals  22.5-23.4 hours of fine-tuning time.
@@
Figure 6: Hyperparameter robustness on extreme years: VITA-Sinusoidal consistently outperforms other baselines across all configurations tested. The CNN-RNN performs considerably worse and is not shown in the plot.
@@
Tables 5 and 9 comprehensively evaluate pretraining benefits across hyperparameter configurations (seed 1234). T-BERT shows consistent but moderate gains: +10.8% mean corn R² (0.647 vs 0.584, $t=9.13$, $p<0.001$) and +5.0% soybean R² (0.693 vs 0.660, $t=4.36$, $p<0.01$), with low variance (±0.014-0.020) indicating stable, initialization-insensitive optimization.
@@
To establish a domain-grounded comparison, we implement a linear regression baseline following the USDA Economic Research Service (ERS) yield modeling framework (Westcott and Jewison 2013). This approach uses agronomically-motivated features (temperature and precipitation during critical growth periods).
@@
Lead Time Performance
Crop VITA Prior 7 Years 5 Years
R² (RMSE) R² (RMSE)
Corn Std. Normal 0.706 (17.1) 0.666 (18.2)
Sinusoidal 0.729 (16.3) 0.697 (17.4)
Soybean Std. Normal 0.698 (5.2) 0.694 (5.3)
Sinusoidal 0.722 (5.0) 0.684 (5.3)
Reduced Temporal Context